OpenStudy (anonymous):

research center survey of 427 women showed that 29.0% of them purchased books online. among the 427 women, what is the number of women who said they purchased books online? find a 95% confidence interval esitmate of the percentage of all women who purchased books online.

OpenStudy (magbak):

Yea you asked a relatively highly leveled question not only stats or A.P stats but collaborative stats and I think no one here is at the level for this question. Sorry. Wish you luck.

OpenStudy (kropot72):

The number of women in the survey who purchased books online is given by: \[427\times0.29=124\ (rounded\ up)\] The survey result forms a binomial distribution where: n = 427 p = 0.29 (1 - p ) = 0.71 The standard distribution is given by: \[s.d.=\sqrt{np(1-p)}=\sqrt{427\times0.29\times0.71}=9.377\] The sample mean is \[np=427\times0.29=123.83\] The confidence interval is estimated from: \[123.83\pm(1.96\times\frac{9.377}{\sqrt{427}})=123.83\pm0.89\] The percentages of women who purchased books online corresponding to these limit are: \[\frac{122.94}{427}\times\frac{100}{1}=28.8\ percent\] and \[\frac{124.72}{427}\times\frac{100}{1}=29.2\ percent\]